OpenStudy (anonymous):

45 = 3b + 69 I have to show my work!!! I get so confused

OpenStudy (anonymous):

problem ?

OpenStudy (anonymous):

not sure which to start with

OpenStudy (anonymous):

what is so confusing ?

OpenStudy (anonymous):

45=72b?

OpenStudy (anonymous):

Simplifying 45 = 3b + 69 Reorder the terms: 45 = 69 + 3b Solving 45 = 69 + 3b

OpenStudy (anonymous):

the whole subject is confusing to me- and my dad cant help me because he doesnt do well in math =(

OpenStudy (anonymous):

you need to find the b ? what's the question ? maybe i can help :)

OpenStudy (anonymous):

3b+69=45 -69 -69 3b=-24 /3 /3 b=-8 there is b and the work if that's what you have to show :)

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

I think all it's asking you is to solve the equation showing your work which is what i have done for you

OpenStudy (anonymous):

you just need to combine like terms .. in your case , in 45=3b+69 , transpose 69 to 45 then it will be 45-69=3b , then perform the operation , -24=3b , divde both sides by three so that the b will remain then b=-8

OpenStudy (anonymous):

so it would be subtract 45 from 69, then divide by 3? and b would be -8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes you could do it either way and the answer for b will still be -8 and both answers are correct

OpenStudy (anonymous):